OpenStudy (anonymous):

How do I find the hole(s) in the Rational Function f(x)=(x-2) / ((x^2)+2x-15) ?

OpenStudy (anonymous):

see if denominator has a factor of (x-2)

OpenStudy (anonymous):

if so, it would cancel the numerator and creates a hole at x=2

OpenStudy (anonymous):

so if it doesn't cancel then there isn't one?

OpenStudy (anonymous):

thats right, if it doesnt cancel no holes

OpenStudy (anonymous):

we will have all vertical asymptotes

OpenStudy (anonymous):

holes occur only when some factor cancels in the denominator, it removes the vertical asymptote and creates a hole

OpenStudy (anonymous):

how do I find a verticle asymptote?

OpenStudy (anonymous):

equate the denominator to 0

OpenStudy (anonymous):

find the solutions by factoring it

OpenStudy (anonymous):

((x^2)+2x-15) = 0

OpenStudy (anonymous):

there can be 2 vertical asymptotes here if nothing cancels the numerator else, 1 vertical asymptote and 1 hole are possible

OpenStudy (anonymous):

((x^2)+2x-15) = 0 (x-3)(x+5)=0 x =3, x = -5 two VAs !

OpenStudy (anonymous):

how do I graph them?

OpenStudy (anonymous):

we show VAs in graph with dotted lines

OpenStudy (anonymous):